# Stabilizing black-box model selection with the inflated argmax

### v1

@@ Abstract
Model selection is the process of choosing from a class of candidate models given data. For instance, methods such as the LASSO and sparse identification of nonlinear dynamics (SINDy) formulate model selection as finding a sparse solution to a linear system of equations determined by training data. However, absent strong assumptions, such methods are highly unstable: if a single data point is removed from the training set, a different model may be selected. This paper presents a new approach to stabilizing model selection that leverages a combination of bagging and an “inflated” argmax operation. Our method selects a small collection of models that all fit the data, and it is stable in that, with high probability, the removal of any training point will result in a collection of selected models that overlaps with the original collection. In addition to developing theoretical guarantees, we illustrate this method in (a) a simulation in which strongly correlated covariates make standard LASSO model selection highly unstable and (b) a Lotka–Volterra model selection problem focused on identifying how competition in an ecosystem influences species’ abundances. In both settings, the proposed method yields stable and compact collections of selected models, outperforming a variety of benchmarks.
@@ 1 Introduction
Scientists often choose between many models to identify key aspects of a given phenomenon. In areas such as economics and biology, models are typically idealized (there is little sense in talking about a ‘true’ model) but also interpretable. The final selection is pragmatically treated as getting at some important truth of the structure of the underlying phenomenon.
A necessary condition for trustworthy model selection procedures is stability, i.e., the selection should not be too sensitive to small changes in the data (Yu,, 2013). With limited information, rich model classes, or complex phenomena, principled methods may not be able to reliably and definitively extract a single best model.
Bagging (Breiman, 1996a, ; Breiman, 1996b, ) has been widely applied to stabilize model selection procedures, see § 3 for examples. When the base procedure is unstable, bagging places near-equal weight on models with similar fits to the data. Bagging can thus express model uncertainty, but, as we show empirically in § 5, selecting the most frequent model among multiple close contenders can still be quite unstable.
@@ 1.1 Our contributions
In this work, we allow model selection procedures to return a set of candidate models, conveying uncertainty about which model is best while returning a single model as often as possible. We formulate a notion of model selection stability based on the idea that our claims to which model is best should remain logically consistent when we drop a small amount of data.
We develop a framework for stabilizing model selection procedures by combining bagged model selection with the inflated argmax (Soloff et al., 2024a, ). We give a theoretical guarantee for model selection stability associated with our framework. We show empirically that the inflated argmax can improve model stability, particularly in the presence of highly correlated covariates, a setting known to yield instabilities in standard model selection algorithms.
2 Our framework
A model selection procedure is a map $\mathcal{M}$ from data $\mathcal{D}$ to a set of selected models. The output $\hat{M}=\mathcal{M}(\mathcal{D})$ is a nonempty subset of all candidate models, denoted $M^{+}$. If $\hat{M}$ is always a singleton, $\mathcal{M}$ is called simple.
@@ 2 Our framework
The model class need not consist of subsets of variables. Each $m\in M^{+}$ could be a different learning rule $f:\mathcal{X}\to\mathcal{Y}$, or a probability distribution $P$ meant to describe the data $\mathcal{D}$. For instance, in graphical model selection, $M^{+}$ is the set of all graphs $G$ on $d$ vertices, where an edge $(j,k)$ in $G$ encodes conditional dependence between variables $j$ and $k$ (Drton and Perlman,, 2004; Friedman et al.,, 2008). In dynamical system identification, $M^{+}$ is a set of possible differential equations governing the dynamics underlying noisy data (Brunton et al.,, 2016). In econometrics, $M^{+}$ could be a collection of competing time series forecasts (Masini et al.,, 2023). To keep our discussion general, the model class is an abstract, countable set $M^{+}$.
We decompose the procedure $\mathcal{M}$ into two stages
$\mathcal{M}=\mathcal{S}\circ\mathcal{A}$ (1)
@@
where $\mathcal{A}$ maps data $\mathcal{D}$ to weights $\hat{w}=\mathcal{A}(\mathcal{D})$ on the model class $M^{+}$, i.e., $\hat{w}_{m}$ is the weight assigned to model $m\in M^{+}$ and is higher for models that ${\cal A}$ deems better fit the data. For instance, in Bayesian model selection, $\hat{w}_{m}$ corresponds to the posterior probability of $m$ given the data $\mathcal{D}$. The selection rule $\mathcal{S}$ takes these weights $\hat{w}$ as input and returns a non-empty set of selected models $\mathcal{S}(\hat{w})\subseteq M^{+}$.
Even if $\mathcal{M}$ is only accessible as a black-box, we can define $\mathcal{S}$ and $\mathcal{A}$ such that (1) holds. Specifically, if
$(\mathcal{A}_{\mathcal{M}}(\mathcal{D}))_{m}:=\frac{1}{|\mathcal{M}(\mathcal{D})|}\mathbbm{1}\{m\in\mathcal{M}(\mathcal{D})\},$ (2)
then we can interpret $\mathcal{M}$ as $\mathop{\!\arg\!\max}\circ\,\mathcal{A}_{\mathcal{M}}$ – i.e., here ${\cal S}=\mathop{\!\arg\!\max}$ is selecting models returned by ${\cal A}$ with the largest weight.
We now provide a brief overview of our pipeline for stabilizing a model weighting algorithm $\mathcal{A}$.
  1. 1.
  2. 2.
    Next, we pass through the bagged weights $\hat{w}$ through the inflated argmax, which selects a set of near-maximizers of $\hat{w}$ that is robust to small perturbations of the weights. We define the inflated argmax in Algorithm 2.
2.1 Bagged model selection
@@
The output of the weighting algorithm $\mathcal{A}$ belongs to the probability simplex $\Delta_{|M^{+}|-1}$. Bagging averages an ensemble of weight vectors fit on different subsets of the data — see Algorithm 1. We denote the resulting bagged algorithm by $\widetilde{\mathcal{A}}_{K,B}$, where $K$ is the number of samples in a bag and $B$ is the number of bags. Sampling with replacement is known as bagging, and sampling without replacement is known as subbagging.
@@
For simple model selection procedures, the ensemble weight $\hat{w}_{m}$ simply counts the fraction of bags wherein model $m$ was selected. Even if the model class $M^{+}$ is countably infinite, the weights $\hat{w}$ will have at most $B$ nonzero entries, so bagging is still tractable.
Algorithm 1 Bagged model selection
@@
Algorithm 2 Inflated Argmax
0:  Weights $\hat{w}$; inflation tolerance $\varepsilon$
  Solve for $c$ such that
$\left(\sum_{m\in M^{+}}(\hat{w}_{m}-c)_{+}\right)^{2}+\sum_{m\in M^{+}}(\hat{w}_{m}-c)^{2}_{+}=\varepsilon^{2}$
  $\left\{m\in M^{+}:\hat{w}_{m}>c+\frac{\varepsilon}{\sqrt{2}}-\sum_{m^{\prime}}(\hat{w}_{m^{\prime}}-c)_{+}\right\}$
In Bayesian inference, $\mathcal{A}(\mathcal{D})$ is a posterior distribution over models $m\in M^{+}$, so $\widetilde{\mathcal{A}}_{K,B}(\mathcal{D})$ is the average of $B$ posteriors fit to subsets of size $K$. This method, known as BayesBag, has been previously applied to Bayesian model selection (Huggins and Miller,, 2023).
Soloff et al., 2024b show that bagged weights are stable, meaning that dropping a single sample cannot change the bagged weights $\hat{w}$ by much in the Euclidean norm. However, in order to translate this into any kind of guarantee on the selected model, a new selection rule $\mathcal{S}$ is needed, as we discuss in the next section.
2.2 The inflated argmax
Soloff et al., 2024a introduced the inflated argmax to guarantee stability in the context of multiclass classification. The inflated argmax, denoted by argmaxε, is a selection rule that will allow us to decide which subset $M$ of $M^{+}$ to return.
@@ 2.2 The inflated argmax
$R_{m}^{\varepsilon}=\left\{w\in\Delta_{|M^{+}|-1}:w_{m}\geq\max_{m^{\prime}\neq m}w_{m^{\prime}}+\frac{\varepsilon}{\sqrt{2}}\right\}.$ (3)
For any $w\in\Delta_{|M^{+}|-1}$ and $\varepsilon>0$, the inflated argmax is defined as
@@
An intuitive explanation of Equation 4 is that $m\in\text{argmax}^{\varepsilon}(w)$ indicates that $m$ would have the largest weight in $w$ by some margin if the weights $w$ were slightly perturbed, where both the size of the margin and the weight perturbations scale with $\varepsilon$. This definition naturally allows for the number of returned models to adapt to the uncertainty in the model selection: for example, if the top two competing models are close in probability, then the inflated argmax will return both models for some choice of small $\varepsilon$. On the other hand, if the top two competing models have a large enough gap in their two weights, the inflated argmax will only return the top model for a choice of small $\varepsilon$. We present a visual representation of the inflated argmax in the supplement.
@@ 2.3 Stability guarantee
In this section, we define stability in the context of model selection and state a theorem guaranteeing the stability of our method.
To assess the stability of $\mathcal{M}$, we measure how the output of $\mathcal{M}$ changes with small perturbations applied to $\mathcal{D}$. We adapt the definition of stability in Soloff et al., 2024a to the context of model selection.
Definition 2.2.
A procedure $\mathcal{M}$ has model selection stability $\delta$ at sample size $n$ if, for all datasets $\mathcal{D}$ with $n$ samples,
In Definition 2.2, $\mathcal{D}^{\setminus i}$ refers to dataset $\mathcal{D}$ with the $i$-th sample removed. In plain language, the inequality in (5) quantifies stability in terms of whether the selected models overlap when we drop a single observation at random.
@@
For any model selection procedure $\mathcal{S}\circ\mathcal{A}$, our method $\textnormal{argmax}^{\varepsilon}\circ\widetilde{A}_{K,B}$ satisfies model selection stability at instability level
@@
This is adapted from and follows immediately from Soloff et al., 2024a (, Theorem 17).
@@
Recall that $\varepsilon$ determines how much we inflate the argmax, so when it’s smaller, the inflated argmax is closer to the standard argmax. Accordingly, when $\varepsilon$ is smaller, Theorem 1 gives a larger $\delta$, reflecting a weaker stability guarantee. Further, note that the instability level $\delta$ decreases (i.e., stability increases) by decreasing the bag size $K$ and increasing the number $B$ of bags.
The size of our model class $|M^{+}|$ has only a small influence on the level of instability of our procedure. For example, when there are two models, $|M^{+}|=2$, the factor $1-\frac{1}{|M^{+}|}=\frac{1}{2}$, giving a constant-factor boost in stability over the case where $M^{+}$ is countably infinite, where $1-\frac{1}{|M^{+}|}$ becomes $1$.
3 Related work
Returning a set of models to acknowledge inherent uncertainty in model selection is an old idea. Early literature on subset selection for linear regression (Spjøtvoll,, 1972) focused on selecting the true model with high probability. More recent works extend the idea of a model confidence set to more general modeling settings (Hansen et al.,, 2011; Zhang et al.,, 2024). These works have a stochastic model for the data and construct a confidence set containing the best-fitting model with high probability. By contrast, our framework places no stochastic assumptions on the data, instead targeting a more modest goal of stability for the selected set.
There is extensive literature on stability in variable selection; see, e.g., Khaire and Dhanalakshmi, (2022) for a recent overview. Work in this area has overwhelmingly focused on “simple” procedures, returning a single subset of variables, and stability is measured by the degree of similarity of the selected subsets of variables. By contrast, Definition 2.2 is not satisfied by simple procedures unless they return the same model most of the time.
Bagging is an important tool for model selection and variable selection (Breiman, 1996b, ). For example, stability selection thresholds the marginal bootstrap inclusion probabilities of each variable (Meinshausen and Bühlmann,, 2010; Shah and Samworth,, 2013). BayesBag (Bühlmann,, 2014) directly averages posterior distributions over different resamplings of the data. Huggins and Miller, (2023) analyze the accuracy and stability properties of the BayesBag approach specifically for Bayesian model selection.
Our work builds on ideas from stable classification. Soloff et al., 2024a introduces a framework for set-valued classification, combining the inflated argmax with bagging to guarantee stability for any base model selection method.
3.1 Alternative selection rules
We discuss the limitations of some alternative approaches to selection $\mathcal{S}$.
The standard argmax.
The default selection rule selects the best model(s) $\hat{M}\in\mathop{\!\arg\!\max}_{m}\hat{w}_{m}$. ($\hat{M}$ may contain more than one model in the case of tied weights.) If the weights contain near-maximizers, the argmax can be sensitive to small perturbations of $\hat{w}$.
Top-$k$.

### v2

@@ Abstract
Model selection is the process of choosing from a class of candidate models given data. For instance, methods such as the LASSO and sparse identification of nonlinear dynamics (SINDy) formulate model selection as finding a sparse solution to a linear system of equations determined by training data. However, absent strong assumptions, such methods are highly unstable: if a single data point is removed from the training set, a different model may be selected. In this paper, we present a new approach to stabilizing model selection with theoretical stability guarantees that leverages a combination of bagging and an “inflated” argmax operation. Our method selects a small collection of models that all fit the data, and it is stable in that, with high probability, the removal of any training point will result in a collection of selected models that overlaps with the original collection. We illustrate this method in (a) a simulation in which strongly correlated covariates make standard LASSO model selection highly unstable, (b) a Lotka–Volterra model selection problem focused on identifying how competition in an ecosystem influences species’ abundances, and (c) a graph subset selection problem using cell-signaling data from proteomics. In these settings, the proposed method yields stable, compact, and accurate collections of selected models, outperforming a variety of benchmarks.
@@ 1 Introduction
Model selection is typically formulated as a procedure to identify the model that “best” represents data from among a set of candidate models. For instance, scientists may choose among many interpretable models to identify salient variables, interactions, or policies, with the goal of developing a fundamental understanding of the natural or social phenomenon underlying experimental data. Examples of model selection include variable selection (choosing a subset of covariates from a larger pool that best explains the response variable or label), selecting a decision tree, selecting the order or memory of an autoregressive model, selecting a kernel function for a support vector machine, selecting the number of clusters in k-means, selecting the number of factors in factor analysis, and more. The issue of model selection is a common problem in a variety of domains, including bioinformatics (Saeys et al.,, 2007), environmental studies (Effrosynidis and Arampatzis,, 2021), and psychology (Vrieze,, 2012).
In this paper, we argue that constraining a procedure to return a single model does not adequately express any uncertainties in the model selection process, and therefore, returning a set of selected model(s) is more appropriate than only ever returning one. An ideal method would return a single model as often as possible while guaranteeing some degree of stability, i.e., the selection should not be too sensitive to small changes in the data (Yu,, 2013).
Especially in model selection scenarios where the goal is to interpret the model to learn something fundamental about the studied process, trustworthiness in the results is paramount. A necessary condition for trustworthy model selection procedures is stability. Conventional methods for model selection provide ad hoc procedures that do not provide theoretical stability guarantees without strong assumptions on the data or model. Our goal is to provide a method that is adaptive to the underlying uncertainty of the model selection process without placing strong assumptions on the data distribution or the models themselves.
In this paper, we propose stabilizing model selection, leveraging ideas from stabilized classification (Soloff et al., 2024b, ), which exhibits desirable theoretical guarantees. We provide comparisons of our proposed approach to widely used conventional model selection methods as well as explore the trade-offs of stability, accuracy, and interpretability of the model selection procedures.
@@ 1.1 Our contributions
In this work, we allow model selection procedures to return a set of candidate models, conveying uncertainty about which model is best while returning a single model as often as possible. We formulate a notion of model selection stability based on the idea that our claims about which model is best should remain logically consistent when we drop a small amount of data.
We develop a framework for stabilizing black-box model selection procedures by combining bagged model selection with the inflated argmax (Soloff et al., 2024b, ), which was formulated in the context of multiclass classification. The key insight of this paper is that if we formulate model selection as a multi-class classification problem, where each class corresponds to one of the candidate models, then we can directly use the inflated argmax and inherit its corresponding theoretical stability guarantees for model selection, which are agnostic to the model class and data distribution. Additionally, the stability guarantee holds even if the “true” model is not considered in the class of candidate models. We emphasize that our approach is highly generic and can be applied to many base model selection tasks, such as variable selection, constructing decision trees, selecting the order of an autoregressive model, and more.
We show empirically that the inflated argmax can improve model selection stability, and for variable selection (a special case of model selection), we similarly show this result for data with correlated covariates, a setting known to yield instabilities in standard model selection algorithms. We compare against and outperform numerous conventional model selection approaches, described in § 2.2. These conventional methods, unlike our approach, are not adaptive to the underlying level of uncertainty of the model selection procedure.
2 Model selection
More formally, model selection focuses on selecting one or more models from among a set of candidate models, denoted $M^{+}$. A model selection procedure is a map $\mathcal{M}$ from data $\mathcal{D}$ to a set of selected models $\hat{M}\subseteq M^{+}$. The output $\hat{M}=\mathcal{M}(\mathcal{D})$ is a nonempty subset of $M^{+}$. If $\mathcal{M}$ always a singleton (i.e., $|\hat{M}|=1$), $\mathcal{M}$ is called simple.
@@ 2 Model selection
The set of all candidate models $M^{+}$ can contain any class of models. For instance, in graphical model selection, $M^{+}$ is the set of all graphs $G$ on $d$ vertices, where an edge $(j,k)$ in $G$ encodes conditional dependence between variables $j$ and $k$ (Drton and Perlman,, 2004; Friedman et al.,, 2008). In dynamical system identification, $M^{+}$ is a set of possible differential equations governing the dynamics underlying noisy data (Brunton et al.,, 2016). In econometrics, $M^{+}$ could be a collection of competing time series forecasts (Masini et al.,, 2023). To keep our discussion general, the model class is an abstract, countable set $M^{+}$.
We decompose the model selection procedure $\mathcal{M}$ into two stages
$\mathcal{M}=\mathcal{S}\circ\mathcal{A},$ (1)
@@
where $\mathcal{A}$ maps data $\mathcal{D}$ to weights $\hat{w}=\mathcal{A}(\mathcal{D})\in\mathbb{R}^{|M^{+}|}$, where for each $m\in M^{+}$, $\hat{w}_{m}$ is the weight assigned to model $m\in M^{+}$, and this weight can be thought of as a probability of each candidate model. The selection rule $\mathcal{S}$ takes these weights $\hat{w}$ as input and returns a non-empty set of selected model(s) $\hat{M}\subseteq M^{+}$ that provide the best fit to the data.
2.1 Stage 1: Assigning weights to candidate models with $\mathcal{A}$
Base algorithm $\mathcal{A}$. A base algorithm $\mathcal{A}$ is any off-the-shelf model weighting algorithm. Many model selection methods lead to weight vectors $\hat{w}$ that place all weight on a single model (i.e., the algorithm $\mathcal{A}$ returns only one model, so $w$ is a vector of zeros with an entry of 1 corresponding to the selected model). Examples of such algorithms are LASSO for variable selection (Tibshirani,, 1996), SINDy for model discovery for dynamical systems (Brunton et al.,, 2016), and graphical LASSO for graph subset selection (Friedman et al.,, 2008), all of which are experimentally explored in this work. A known issue of these procedures is that these algorithms can produce unstable model selections. One mechanism that has been shown to help its stability is bagging.
@@ 2.1 Stage 1: Assigning weights to candidate models with $\mathcal{A}$
Bagged algorithm $\tilde{\mathcal{A}}_{K,B}$. We denote bagging a base algorithm $\mathcal{A}$ as $\tilde{\mathcal{A}}_{K,B}$, where $K$ is the number of samples per bag, and $B$ is the number of bags. The output of the weighting algorithm $\tilde{\mathcal{A}}_{K,B}$ must be bounded. For simplicity, we assume that the weights can be normalized such that $\hat{w}$ belongs to the probability simplex $\Delta_{|M^{+}|-1}$.  [^1] The probability simplex $\Delta_{|M^{+}|-1}$ consists of nonnegative weights $w$ that sum to one: $\sum_{m\in M^{+}}w_{m}=1$. Bagging averages an ensemble of weight vectors fit on different randomly sampled bags — see Algorithm 1. Sampling with replacement is known as bagging, and sampling without replacement is known as subbagging.
@@
For simple model selection procedures, the ensemble weight $\hat{w}_{m}$ simply counts the fraction of bags wherein model $m$ was selected. Even if the set of candidate models is infinite (i.e., $|M^{+}|=\infty$), the weights $\hat{w}$ will have at most $B$ nonzero entries, so bagging is still tractable. Generally, increasing the number of bags $B$ used in $\tilde{\mathcal{A}}_{K,B}$ provides a more stability, as we empirically show in Figures 12 and 13.
Algorithm 1 Bagged model selection (Breiman, 1996a, ; Breiman, 1996b, )
@@
2.2 Stage 2: Conventional selection rules $\mathcal{S}$
We discuss conventional selection approches $\mathcal{S}$ and their limitations.
The standard argmax. The default selection rule selects the best model(s) $\hat{M}\in\mathop{\!\arg\!\max}_{m}\hat{w}_{m}$. ($\hat{M}$ may contain more than one model in the case of exact ties for the top weight.) If the weights contain near-maximizers, the argmax can be sensitive to small perturbations of $\hat{w}$, meaning that a different model may be selected with a small perturbation to the training data.
[…]
Stability selection. In the context of variable selection, the stability selection method (Meinshausen and Bühlmann,, 2010) selects a single model based on whether each covariate is included in the model with probability at least $\tau$. Concretely, consider a vector of weights $\hat{w}$ indexed by $M^{+}=2^{[d]}$, where $\hat{w}_{m}$ roughly represents the probability that $m$ is the correct model. Define the marginal variable probabilities $\hat{\Pi}_{j}=\sum_{\begin{subarray}{c}m\subseteq[d]:j\in m\end{subarray}}\hat{w}_{m},$ for $j=1,\ldots,d,$ where $j\in m$ corresponds to variable $j$ being included in model $m$. Stability selection returns
$\text{ip}_{\tau}(\hat{w}):=\Big{\{}j\in[d]:\hat{\Pi}_{j}\geq\tau\Big{\}},$
for some use-specified threshold $\tau$. Similar to the standard argmax (without ties), $\text{ip}_{\tau}(\hat{w})$ only returns a collection of variables, which together constitute a single model. This selection procedure can be unstable if some marginal probabilities $\hat{\Pi}_{j}$ are near the threshold $\tau$. It also may fail to adapt to relationships among variables (see Appendix A for an example).
2.3 Model selection stability
To assess the stability of $\mathcal{M}$, we measure how the output of $\mathcal{M}$ changes with small perturbations to $\mathcal{D}$. We adapt the definition of stability in Soloff et al., 2024b to the context of model selection.
Definition 2.1.
A procedure $\mathcal{M}$ has model selection stability $1-\delta$ at sample size $n$ if, for all datasets $\mathcal{D}$ with $n$ samples,
In Definition 2.1, $\mathcal{D}^{\setminus i}$ refers to dataset $\mathcal{D}$ with the $i$-th sample removed. In plain language, the inequality in (2) quantifies stability in terms of whether the sets of selected models overlap when we drop a single observation at random. Note that this notion of model selection stability holds for any “black box” model selection method of the form (1), in contrast to prior works that focus exclusively on variable selection (e.g., Meinshausen and Bühlmann, (2010)).
3 Related work
Returning a model selection set. Returning a set of models to acknowledge inherent uncertainty in model selection is an old idea. Early literature on subset selection for linear regression (Spjøtvoll,, 1972) focused on selecting the true model with high probability. More recent works extend the idea of a model confidence set to more general modeling settings (Hansen et al.,, 2011; Zhang et al.,, 2024). These works have a stochastic model for the data and construct a confidence set containing the best-fitting model with high probability. By contrast, our framework places no stochastic assumptions on the data, instead targeting a more modest goal of stability for the selected set.
Stability in variable selection. There is extensive literature on stability in variable selection; see, e.g., Khaire and Dhanalakshmi, (2022) for a recent overview. Work in this area has overwhelmingly focused on “simple” procedures, returning a single subset of variables, and stability is measured by the degree of similarity of the selected subsets of variables. By contrast, Definition 2.1 is not satisfied by simple procedures unless they return the same model most of the time.
Bagging to improve stability. Bagging is an important tool for model selection and variable selection, and has been widely applied to stabilize model selection procedures (Breiman, 1996a, ; Breiman, 1996b, ). For example, stability selection thresholds the marginal bootstrap inclusion probabilities of each variable (Meinshausen and Bühlmann,, 2010; Shah and Samworth,, 2013). Additionally, BayesBag (Bühlmann,, 2014) directly averages posterior distributions over different resamplings of the data. Huggins and Miller, (2023) analyze the accuracy and stability properties of the BayesBag approach specifically for Bayesian model selection. Soloff et al., 2024a ; Soloff et al., 2024c shows that bagged weights are stable, meaning that dropping a single sample cannot change the bagged weights $\hat{w}$ by much in the Euclidean norm. Bagging can thus express model uncertainty, but, as we show empirically in § 5, selecting the most frequent model among multiple close contenders can still be quite unstable. In order to achieve any stability guarantees of the selected model(s), a new selection rule  $\mathcal{S}$ is needed, as discussed in § 4.1.
Stabilizing classification. Our work builds on ideas from stable classification. Soloff et al., 2024b introduces a framework for set-valued classification, combining the inflated argmax with bagging to guarantee stability for any base model selection method.
4 Our approach
We now provide a brief overview of our pipeline for stabilizing a model selection procedure $\mathcal{M}={\cal S}\circ{\cal A}$.
  1. 1.
  2. 2.
    Next, we pass the bagged weights $\hat{w}$ through the inflated argmax, which selects a set of near-maximizers of $\hat{w}$ that is robust to small perturbations of the weights. We define the inflated argmax for model selection in Equation 4.
4.1 The inflated argmax for model selection
In this section, we propose an alternative selection procedure $\mathcal{S}$ based on the inflated argmax. Soloff et al., 2024b introduced the inflated argmax to guarantee stability in the context of multiclass classification. A key insight of this paper is that if we formulate model selection as a multi-class classification problem, where each class corresponds to one of the candidate models, then we can directly use the inflated argmax and inherit its corresponding theoretical stability guarantees for model selection. With this insight, the inflated argmax, an alternative model selection criteria $\mathcal{S}$ that we denotd by argmaxε, will allow us to decide which subset $\hat{M}$ of $M^{+}$ to return.
@@ 4.1 The inflated argmax for model selection
$R_{m}^{\varepsilon}=\left\{w\in\Delta_{|M^{+}|-1}:w_{m}\geq\max_{m^{\prime}\neq m}w_{m^{\prime}}+\frac{\varepsilon}{\sqrt{2}}\right\},$ (3)
where $\Delta_{|M^{+}|-1}$ is a probability simplex. For any $w\in\Delta_{|M^{+}|-1}$ and $\varepsilon>0$, the inflated argmax is defined as
@@
A visualization of a simple example of argmaxϵ where more than one model is returned is shown in Figure 1. An intuitive explanation of Equation 4 is that $m\in\text{argmax}^{\varepsilon}(w)$ indicates that $m$ would have the largest weight in $w$ by some margin if the weights $w$ were slightly perturbed, where both the size of the margin and the weight perturbations scale with $\varepsilon$. This definition naturally allows for the number of returned models to adapt to the uncertainty in the model selection: for example, if the top two competing models are close in probability, then the inflated argmax will return both models for some choice of small $\varepsilon$. On the other hand, if the top two competing models have a large enough gap in their two weights, the inflated argmax will only return the top model for a choice of small $\varepsilon$.
The parameter $\varepsilon$ is not something to be tuned. It is chosen to satisfy a user-specified stability tolerance; see § 4.2.
Figure 1: Visualization of argmaxε when $|M^{+}|=3$. The weights assigned to each model $\mathcal{A}$ can be mapped onto a 3-dimensional probability simplex as shown. The example weight vector $w=\left[\frac{1}{9},\frac{4}{9},\frac{4}{9}\right]$ lands in an area that is within $\varepsilon$ distance of regions $R_{2}^{\varepsilon}$ and $R_{3}^{\varepsilon}$, which are the regions corresponding to returning models $m_{2}$ and $m_{3}$, respectively. Because this weight vector is in an “uncertain” region based on $\varepsilon$ for distinguishing whether model $m_{2}$ or $m_{3}$ provides a better fit to the data, argmaxε will return both models: $\{m_{2},m_{3}\}$. In contrast, weight vector $w=[\frac{1}{9},\frac{1}{9},\frac{7}{9}]$ would land in the area labeled $R_{3}^{\varepsilon}$, and argmaxε will return $m_{3}$ only.
@@ 4.2 Stability guarantee
In this section, state a theorem guaranteeing the stability of our approach based on our definition of stability in Definition 2.1.
@@
(Adapted from Soloff et al., 2024b (, Theorem 17).) For any model selection procedure $\mathcal{S}\circ\mathcal{A}$, our method $\textnormal{argmax}^{\varepsilon}\circ\widetilde{A}_{K,B}$ satisfies model selection stability at instability level
@@
The main use for this theorem is that the user can choose a tolerable worst-case instability $\delta$ for their model selection, input the known parameters of the problem (i.e., $|M^{+}|$ candidate models, $K$ samples per bag, $n$ total samples, and $B$ bags), and directly solve for the $\varepsilon$ that should be used for argmaxε.
@@
Recall that $\varepsilon$ determines how much we inflate the argmax, so when it’s smaller, the inflated argmax is closer to the standard argmax. Accordingly, when $\varepsilon$ is smaller, Theorem 1 gives a larger $\delta$, reflecting a weaker stability guarantee. Further, note that the instability level $\delta$ decreases (i.e., stability increases) by decreasing the bag size $K$ and increasing the number $B$ of bags. The size of our model class $|M^{+}|$ has only a small influence on the level of instability of our procedure. For example, when there are two models, $|M^{+}|=2$, the factor $1-\frac{1}{|M^{+}|}=\frac{1}{2}$, giving a constant-factor boost in stability over the case where $M^{+}$ is countably infinite, where $1-\frac{1}{|M^{+}|}$ becomes $1$.
This theoretical result is the first to provide theoretical guarantees for model selection stability that is agnostic to the underlying models, base algorithm, and data distribution.
5 Experiments
Our experiments assess and compare the stability of five different model selection approaches, described in Table 1.
NOTATION SELECTION CRITERION
We perform experiments in three settings: identifying important variables in a synthetic dataset that contains a small collection of highly correlated variables (see Appendix C), identifying the latent Lotka-Volterra governing differential equations from noisy trajectory data generated from these equations, and constructing a graph to represent cell-signaling pathways of 11 proteins. In the first two settings, we generate synthetic datasets for $N$ trials, each contaminated with independent Gaussian noise, allowing us to examine the stability across a collection of datasets where we know the ground truth $m^{*}$.
In each experiment, we compute weights $\hat{w}$ across the set candidate models $M^{+}$ using a base algorithm $\mathcal{A}$ and a subbagged version of $\mathcal{A}$ denoted $\tilde{\mathcal{A}}_{K,B}$. Based on (2), we compute the stability for trial $j\in[N]$ using
This definition of stability measures, for trial $j$, the proportion of LOO selected models $\hat{M}_{j}^{\setminus i}$ for $i\in[n]$ that have no overlap with the set of selected models $\hat{M}_{j}$ returned by training on the entire dataset. We also compute the empirical cumulative distribution function (CDF)
$\frac{1}{N}\sum_{j=1}^{N}\mathbbm{1}\{\delta_{j}\leq\delta\},$ (6)
as a function of $\delta\in[0,1]$ to highlight variation across trials with different random seeds (shown in Figure 3 and Figure 7). Curves higher on the plot are better in the sense that the corresponding method achieves model selection stability $\delta$ (Definition 2.1) for a larger fraction of trials, and our primary interest is in small $\delta$ (i.e., high stability).
$\frac{1}{N}\sum_{j=1}^{N}\frac{\mathbbm{1}\{m^{*}\in\hat{M}_{j}\}}{|\hat{M}_{j}|},$ (7)
where $m^{*}$ is the “true” data generating model in the synthetic experiments. This measure down-weights the accuracy by the number of models returned $|\hat{M}_{j}|$. This metric is used to compare methods in the center column of Figure 3.
5.1 Model discovery of Lotka-Volterra dynamics
$\dot{u}^{(1)}(t)$ $=\;\;\,\alpha u^{(1)}(t)-\beta u^{(1)}(t)u^{(2)}(t),$ (8a)
$\dot{u}^{(2)}(t)$ $=-\gamma u^{(2)}(t)+\zeta u^{(1)}(t)u^{(2)}(t),$ (8b)
where $u^{(1)}$ is the population density of prey, $u^{(2)}$ is the population density of predators, $\dot{u}^{(1)}$ and $\dot{u}^{(2)}$ are the population growth rates, and $t$ corresponds to time. The parameters $\alpha,\beta,\gamma,\zeta>0$ respectively refer to the prey’s growth rate, the predator’s impact on the prey’s death rate, the predator’s death rate, and the prey’s impact on the predator’s growth rate (Lotka,, 1925). We now imagine we do not know our system is governed by the Lotka-Volterra equations and wish to discover the governing model from data. Our concern is in recovering the correct terms rather than estimating the parameters.
5.1.1 Base algorithm $\mathcal{A}$: Sparse Identification of Nonlinear Dynamical Systems (SINDy)
One application of sparse regression is the discovery of governing equations of a system from observations. Specifically, we seek to learn the governing equations of a system of the form
$\frac{d}{dt}u(t)=f(u(t)),$ (9)
where $u(t)=\begin{bmatrix}u^{(1)}(t),u^{(2)}(t),\cdots,u^{(d)}(t)\end{bmatrix}^{\top}$ is a $d$-dimensional state vector containing quantities of interest at time $t$, and $f$ is the function describing how the derivative of $u(t)$ evolves with time. The Sparse Identification of Nonlinear Dynamical Systems (SINDy, Brunton et al., (2016)) algorithm represents a differential equation of the form (9) as a weighted sum of terms in a library of functional forms,
$\dot{u}(t)=\Theta(u)\Xi,$ (10)
where $\Theta(u)\in\mathbb{R}^{n\times p}$ ($n$ samples of $p$ functionals) is the library of functionals of $u(t)$ (e.g., polynomial transformations), and $\Xi\in\mathbb{R}^{p\times d}$ are the fitted coefficients. $\Xi$ is assumed to be sparse, so our goal is to find which elements of $\Xi$ should be nonzero, which will immediately give the estimated structure of the governing equations. SINDy finds a sparse solution to $\Xi$ by solving a sparse linear regression problem; see Appendix B for more details on sparse regression.